# Global existence and boundedness in an attraction-repulsion chemotaxis system with nonlocal logistic source and sublinear productions

### v1

@@ 2 Preliminaries and local existence
$+C_{2}\int^{t}_{0}e^{-\alpha(t-s)}\left(1+\left(t-s\right)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{p}-\frac{1}{q_{0}\sigma}\right)}\right)e^{-\lambda(t-s)}\Big{\|}{u}^{l}(\cdot,s)\Big{\|}_{{\mathcal{L}^{p}}(\Omega)}\mathrm{d}s,$ (2.7)
where $C_{1}$ and $C_{2}$ are positive constants. Here, we can guarantee that $\frac{1}{2}+\frac{n}{2}\left(\frac{1}{p}-\frac{1}{q_{0}\sigma}\right)<1$, because of our assumption $q_{0}\sigma<\frac{np}{(n-p)}$ and $q_{0}\sigma\leq q$. Now, adopting the Gamma function, we ensure $\int_{0}^{\infty}e^{-\alpha\psi}\left(1+\psi^{-\frac{1}{2}-\frac{n}{2}(\frac{1}{p}-\frac{1}{q_{0}\sigma})}\right)e^{-\lambda\psi}<\infty$. Using the Interpolation inequality with Lemma 3.2, to get
$\big{\|}{u}^{l}(\cdot,t)\big{\|}_{{\mathcal{L}^{p}}(\Omega)}\leq|\Omega|^{\frac{1-l}{p}}\left(\int_{\Omega}u(\cdot,t)^{p}\right)^{\frac{l}{p}}\leq C_{3}\|{u}(\cdot,t)\big{\|}^{l}_{{\mathcal{L}^{p}}(\Omega)},$ (2.8)
where $l\in(0,1)$. Substituting (2.8) in to (2.7), we arrive at
$\big{\|}\nabla{v}(\cdot,t)\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$ $\leq C_{1}\|{v}_{0}\|_{{\mathcal{W}^{1,q}}(\Omega)}+C_{4},$
where $C_{4}$ is positive constant. Finally, we obtain
$\big{\|}\nabla{v}(\cdot,t)\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$ $\leq C_{5},\qquad\qquad\forall\,t\in(s_{0},T_{\mathrm{max}}).$ (2.9)
@@
$\big{\|}{\nabla{w}}(\cdot,t)\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$ $\leq C_{6},\qquad\qquad\forall\,t\in(s_{0},T_{\mathrm{max}}).$ (2.10)
where $C_{5}$ and $C_{6}>0$. We can attain the following representation by adopting the variation of constants formula
@@
$+\mu\int^{t}_{0}e^{d_{1}(t-s)\Delta}{u}^{m}(\cdot,s)\Big{(}1-\int_{\Omega}{u}(\cdot,s)\mathrm{d}x\Big{)}\mathrm{d}s.$ (2.11)
@@
$\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}\leq$ $\>\Big{\|}e^{d_{1}t\Delta}{u}(\cdot,0)\Big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}+\chi\int^{t}_{0}\Big{\|}e^{d_{1}(t-s)\Delta}\nabla\cdot\Big{(}{u}^{k}(\cdot,s){\nabla{v}(\cdot,s)}\Big{)}\Big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}\mathrm{d}s$
$+\xi\int^{t}_{0}\Big{\|}e^{d_{1}(t-s)\Delta}\nabla\cdot\Big{(}{u}^{k}(\cdot,s){\nabla{w}(\cdot,s)}\Big{)}\Big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}\mathrm{d}s$
$+\mu\int^{t}_{0}\left\|e^{d_{1}(t-s)\Delta}{u}^{m}(\cdot,s)\left(1-M(t)\right)\right\|_{{\mathcal{L}^{\infty}}(\Omega)}\mathrm{d}s,$
for all $t\in(0,T_{\mathrm{max}})$. Now, applying the Neumann heat semigroup [42] with $C_{7}$ and $C_{8}>0$ satisfying
@@
$+C_{7}\int^{t}_{0}\left(1+\left(t-s\right)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q_{0}}-\frac{1}{\infty}\right)}\right)e^{-\lambda(t-s)}\Big{\|}{u}^{k}(\cdot,s){\nabla{v}(\cdot,s)}\Big{\|}_{{\mathcal{L}^{q_{0}}}(\Omega)}\mathrm{d}s$
$+C_{8}\int^{t}_{0}\left(1+\left(t-s\right)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q_{0}}-\frac{1}{\infty}\right)}\right)e^{-\lambda(t-s)}\Big{\|}{u}^{k}(\cdot,s){\nabla{w}(\cdot,s)}\Big{\|}_{{\mathcal{L}^{q_{0}}}(\Omega)}\mathrm{d}s$
$+|1-M_{0}|C_{9}\int^{t}_{0}\left(1+\left(t-s\right)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q_{0}}-\frac{1}{\infty}\right)}\right)e^{-\lambda(t-s)}\Big{\|}{u}^{m}(\cdot,s)\Big{\|}_{{\mathcal{L}^{q_{0}}}(\Omega)}\mathrm{d}s,$ (2.12)
@@
$\leq\big{\|}{u}(\cdot,s)\big{\|}^{k}_{L^{q_{0}\widehat{\sigma}k}(\Omega)}\;\big{\|}{\nabla{v}(\cdot,s)}\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$
@@
$\leq\Big{\|}{u}(\cdot,s)\Big{\|}_{{\mathcal{L}^{1}}(\Omega)}^{\frac{1}{q_{0}\widehat{\sigma}}}\;\Big{\|}{u}(\cdot,s)\Big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}^{k-\frac{1}{q_{0}\widehat{\sigma}}}\;\big{\|}{\nabla{v}(\cdot,s)}\big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$
$\leq C_{10}\Big{\|}{u}(\cdot,s)\Big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}^{k-\frac{1}{q_{0}\widehat{\sigma}}}.$
Now, using the Young’s inequality with $q_{0}\widehat{\sigma}>\frac{n}{2}$, gives
$\big{\|}{u}^{k}(\cdot,s){\nabla{v}(\cdot,s)}\big{\|}_{L^{q_{0}}(\Omega)}$ $\leq\frac{1}{6}\big{\|}{u}(\cdot,s)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}+C_{11}.$ (2.13)
Similarly, we can attain
$\big{\|}{u}^{k}(\cdot,s){\nabla{w}(\cdot,s)}\big{\|}_{L^{q_{0}}(\Omega)}$ $\leq\frac{1}{6}\big{\|}{u}(\cdot,s)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}+C_{12},$ (2.14)
where $\widehat{\sigma}$ is the dual exponent of $\sigma$ and ${\zeta_{2}=\frac{1}{q_{0}\widehat{\sigma}k}\in(0,1)}$, for all $s\in(s_{0},T_{\mathrm{max}})$. Also,
$\big{\|}{u}^{m}(\cdot,s)\big{\|}_{{\mathcal{L}^{q_{0}}}(\Omega)}=$ $\Big{\|}{u}(\cdot,s)\Big{\|}^{m}_{L^{q_{0}m}(\Omega)}$
@@
where ${\zeta_{3}=\frac{1}{q_{0}m}\in(0,1)}$, for all $s\in(s_{0},T_{\mathrm{max}})$ and $C_{13}>0$. Once again the Young’s inequality with $q_{0}>\frac{n}{2}$, gives
$\big{\|}{u}^{m}(\cdot,s)\big{\|}_{{\mathcal{L}^{q_{0}}}(\Omega)}\leq$ $C_{13}\Big{\|}{u}(\cdot,s)\Big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}^{m-\frac{1}{q_{0}}}\leq\frac{1}{6}\big{\|}{u}(\cdot,s)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}+C_{14}.$ (2.15)
Inserting (2.13)-(2.15) in (2.12), it follows that
$\sup_{t\in(s_{0},T)}\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}\leq\frac{1}{2}\sup_{t\in(s_{0},T)}\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}+C_{15},$
for all $T\in(s_{0},T_{\mathrm{max}})$. Finally, we conclude that
$\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}\leq C_{16},\qquad\qquad\qquad\forall\,t\in(s_{0},T_{\mathrm{max}}).$ (2.16)
where the constant $C_{16}$ is positive. ∎
@@ 3 Global existence of classical solutions
In this section, we prove the global existence and boundedness of the solution to (1.6). First we derive ${\mathcal{L}^{p}}(\Omega)$ norm for ${u}$. Given any $s_{0}\in(0,T_{\mathrm{max}})$ such that $s_{0}<1$, Lemma 2.1 gives ${u}(\cdot,s_{0}),{v}(\cdot,s_{0}),{w}(\cdot,s_{0})\in{\mathcal{C}^{2}(\overline{\Omega})}$ with $\frac{\partial{v}(\cdot,s_{0})}{\partial\nu}=0$ and $\frac{\partial{w}(\cdot,s_{0})}{\partial\nu}=0$, we pick $C>0$ such that
$\left\{\begin{array}[]{rrll}&&\sup\limits_{0\leq s\leq s_{0}}\|{u}(\cdot,s)\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C,\quad\sup\limits_{0\leq s\leq s_{0}}\|{v}(\cdot,s)\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C,\quad\sup\limits_{0\leq s\leq s_{0}}\|{w}(\cdot,s)\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C,\\ \\ &&\|\Delta{v}(\cdot,s_{0})\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C\quad\mathrm{and}\quad\|\Delta{w}(\cdot,s_{0})\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C.\end{array}\right.$ (3.4)
@@
Suppose that $\Omega\subset\mathbb{R}^{n}$, $n\geq 1$ is an open, bounded domain with smooth boundary. Assume that $p+m-1>p^{\prime}>1$ and $M_{0}<1$. If $m$ take the values,
11 11 $<m<1+\frac{2}{n},\hskip 56.9055pt\mathrm{if}\quad n\geq 3,$
11 11 $<m<2,\hskip 79.6678pt\mathrm{if}\quad n=1,2,$
then we have
$\int_{\Omega}(1-M_{0}){{u}}^{p+m-1}$ $\leq\frac{2d_{1}(p-1)}{p^{2}\mu}\int_{\Omega}\big{|}{\nabla{u}}^{\frac{p}{2}}\big{|}^{2}+C,$ (3.5)
@@
Using the Hölder’s inequality
$\int_{\Omega}{{u}}^{p+m-1}=$ $\int_{\Omega}{u}^{\frac{p}{2}\frac{2\gamma(p+m-1)}{p}}{u}^{\frac{p}{2}\frac{2(1-\gamma)(p+m-1)}{p}}$
$\leq$ $\left(\int_{\Omega}\left({u}^{\frac{p}{2}}\right)^{\frac{2\gamma(p+m-1)}{p}\frac{rp}{2\gamma(p+m-1)}}\right)^{\frac{2\gamma(p+m-1)}{rp}}$
×$\times\left(\int_{\Omega}\left({u}^{\frac{p}{2}}\right)^{\frac{2(1-\gamma)(p+m-1)}{p}\frac{p}{2(1-\gamma)(p+m-1)}\frac{2p^{\prime}}{p}}\right)^{\frac{2(1-\gamma)(p+m-1)}{p}\frac{p}{2p^{\prime}}}$
$\leq$ $\left(\int_{\Omega}\left({u}^{\frac{p}{2}}\right)^{r}\right)^{\frac{1}{r}\frac{2\gamma(p+m-1)}{p}}\left(\int_{\Omega}\left({u}^{\frac{p}{2}}\right)^{\frac{2p^{\prime}}{p}}\right)^{\frac{p}{2p^{\prime}}\frac{2(1-\gamma)(p+m-1)}{p}},$
from the above, we obtain the estimate
$\int_{\Omega}{{u}}^{p+m-1}\leq$ $\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{\frac{2\gamma(p+m-1)}{p}}_{{\mathcal{L}^{r}}(\Omega)}\Big{\|}{u}^{\frac{p}{2}}\Big{\|}^{\frac{2(1-\gamma)(p+m-1)}{p}}_{{\mathcal{L}^{\frac{2p^{\prime}}{p}}}(\Omega)}.$ (3.6)
Using the Sobolev embedding theorem, we get the estimate from (3.6)
@@
Where $\gamma$ is the exponent from the Hölder’s inequality, such that
$\frac{2\gamma(p+m-1)}{rp}+\frac{(1-\gamma)(p+m-1)}{p^{\prime}}$ $=1$
$\frac{2\gamma}{rp}+\frac{(1-\gamma)}{p^{\prime}}$ $=\frac{1}{p+m-1}$
$\frac{\gamma}{r}+\frac{p}{2p^{\prime}}-\frac{\gamma p}{2p^{\prime}}$ $=\frac{p}{2(p+m-1)}$
$\gamma\left[\frac{1}{r}-\frac{p}{2p^{\prime}}\right]$ $=\frac{p}{2(p+m-1)}-\frac{p}{2p^{\prime}},$
we have,
and $r$ satisfies
$r=\left\{\begin{array}[]{rrll}&&\frac{2n}{n-2},&\quad\mbox{if}\quad n\geq 3,\\ \\ &&\left(\frac{2(p+m-1)}{p},\infty\right),&\quad\mbox{if}\quad n=2,\\ \\ &&\infty,&\quad\mbox{if}\quad n=1.\end{array}\right.$
@@
where $a=\frac{2(1-\gamma)(p+m-1)}{p}\frac{1}{1-\frac{\gamma(p+m-1)}{p}}$. Take $p^{\prime}>\frac{n(m-1)}{2}$ with simple computation, we ensure that
@@
In addition, we have the following estimates

### v2

@@ 2 Preliminaries and local existence
$+C_{2}\int^{t}_{s_{0}}e^{-\alpha(t-s)}\left(1+\left(t-s\right)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{p}-\frac{1}{q_{0}\sigma}\right)}\right)e^{-d_{2}\lambda(t-s)}\big\|{u}^{l}(\cdot,s)\big\|_{{\mathcal{L}^{p}}(\Omega)}\mathrm{d}s,$ (2.2)
where $C_{1}$ and $C_{2}$ are positive constants. Here, we can guarantee that $\frac{1}{2}+\frac{n}{2}\left(\frac{1}{p}-\frac{1}{q_{0}\sigma}\right)<1$, because of our assumption $q_{0}\sigma<\frac{np}{(n-p)}$. Now, adopting the Gamma function, we ensure $\int_{0}^{\infty}e^{-\alpha\psi}\left(1+\psi^{-\frac{1}{2}-\frac{n}{2}(\frac{1}{p}-\frac{1}{q_{0}\sigma})}\right)e^{-d_{2}\lambda\psi}<\infty$. Using the Hölder’s inequality, to get
$\big\|{u}^{l}(\cdot,t)\big\|_{{\mathcal{L}^{p}}(\Omega)}=\left(\int_{\Omega}u^{lp}(\cdot,t)\right)^{\frac{l}{p}}\leq|\Omega|^{\frac{1-l}{p}}\left(\int_{\Omega}u^{p}(\cdot,t)\right)^{\frac{l}{p}}\leq C_{3}\|{u}(\cdot,t)\big\|^{l}_{{\mathcal{L}^{p}}(\Omega)},\quad\forall t\in(s_{0},T_{\mathrm{max}}),$ (2.3)
where $l\in(0,1)$ and $C_{3}>0$. Substituting (2.3) in to (2.2), we arrive at
$\big\|\nabla{v}(\cdot,t)\big\|_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$ $\leq C_{1}\|{v}_{0}\|_{{\mathcal{W}^{1,q}}(\Omega)}+C_{4},\qquad\forall t\in(s_{0},T_{\mathrm{max}})$
where $C_{4}$ is positive constant. Finally, we obtain with the help of Lemma 2.1 and $s_{0}<1$
$\big\|\nabla{v}(\cdot,t)\big\|_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$ $\leq C_{5},\qquad\qquad\forall\,t\in(0,T_{\mathrm{max}}).$ (2.4)
@@
$\big\|{\nabla{w}}(\cdot,t)\big\|_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$ $\leq C_{6},\qquad\qquad\forall\,t\in(0,T_{\mathrm{max}}).$ (2.5)
where $C_{5}$ and $C_{6}>0$. Set $t\in(0,T_{\mathrm{max}})$ and we attain the following representation by adopting the variation of constants formula
@@
$+\mu\int^{t}_{0}e^{d_{1}(t-s)\Delta}{u}^{m}(\cdot,s)\Big(1-\int_{\Omega}{u}(\cdot,s)\Big)\mathrm{d}s.$
@@
$\big\|{u}(\cdot,t)\big\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq$ $\>\Big\|e^{d_{1}t\Delta}{u}(\cdot,0)\Big\|_{{\mathcal{L}^{\infty}}(\Omega)}+\chi\int^{t}_{0}\Big\|\nabla e^{d_{1}(t-s)\Delta}\Big({u}^{k}(\cdot,s){\nabla{v}(\cdot,s)}\Big)\Big\|_{{\mathcal{L}^{\infty}}(\Omega)}\mathrm{d}s$
$+\xi\int^{t}_{0}\Big\|\nabla e^{d_{1}(t-s)\Delta}\Big({u}^{k}(\cdot,s){\nabla{w}(\cdot,s)}\Big)\Big\|_{{\mathcal{L}^{\infty}}(\Omega)}\mathrm{d}s$
$+\mu\int^{t}_{0}\left\|e^{d_{1}(t-s)\Delta}{u}^{m}(\cdot,s)\big(1-M(s)\big)\right\|_{{\mathcal{L}^{\infty}}(\Omega)}\mathrm{d}s,$
for all $t\in(0,T_{\mathrm{max}})$. Now, applying the Neumann heat semigroup [43] with $C_{7},C_{8}$ and $C_{9}>0$ satisfying
@@
$+C_{7}\int^{t}_{0}\left(1+\left(t-s\right)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q_{0}}-\frac{1}{\infty}\right)}\right)e^{-d_{1}\lambda(t-s)}\Big\|{u}^{k}(\cdot,s){\nabla{v}(\cdot,s)}\Big\|_{{\mathcal{L}^{q_{0}}}(\Omega)}\mathrm{d}s$
$+C_{8}\int^{t}_{0}\left(1+\left(t-s\right)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q_{0}}-\frac{1}{\infty}\right)}\right)e^{-d_{1}\lambda(t-s)}\Big\|{u}^{k}(\cdot,s){\nabla{w}(\cdot,s)}\Big\|_{{\mathcal{L}^{q_{0}}}(\Omega)}\mathrm{d}s$
$+|1-M_{0}|C_{9}\int^{t}_{0}\left(1+\left(t-s\right)^{-\frac{n}{2}\left(\frac{1}{q_{0}}-\frac{1}{\infty}\right)}\right)e^{-d_{1}\lambda(t-s)}\Big\|{u}^{m}(\cdot,s)\Big\|_{{\mathcal{L}^{q_{0}}}(\Omega)}\mathrm{d}s,$ (2.6)
@@
$=\big\|{u}(\cdot,s)\big\|^{k}_{L^{q_{0}\widehat{\sigma}k}(\Omega)}\;\big\|{\nabla{v}(\cdot,s)}\big\|_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$
@@
$=\Big\|{u}(\cdot,s)\Big\|_{{\mathcal{L}^{1}}(\Omega)}^{\frac{1}{q_{0}\widehat{\sigma}}}\;\Big\|{u}(\cdot,s)\Big\|_{{\mathcal{L}^{\infty}}(\Omega)}^{k-\frac{1}{q_{0}\widehat{\sigma}}}\;\big\|{\nabla{v}(\cdot,s)}\big\|_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}$
$\leq C_{10}\Big\|{u}(\cdot,s)\Big\|_{{\mathcal{L}^{\infty}}(\Omega)}^{k-\frac{1}{q_{0}\widehat{\sigma}}}$,
where $C_{10}>0$. Now, using the Young’s inequality with $q_{0}\widehat{\sigma}>\frac{n}{2}$ and $C_{11}>0$, gives
$\big\|{u}^{k}(\cdot,s){\nabla{v}(\cdot,s)}\big\|_{{\mathcal{L}^{q_{0}}}(\Omega)}$ $\leq\frac{1}{6}\big\|{u}(\cdot,s)\big\|_{{\mathcal{L}^{\infty}}(\Omega)}+C_{11}.$ (2.7)
Similarly, we can attain with $C_{12}>0$
$\big\|{u}^{k}(\cdot,s){\nabla{w}(\cdot,s)}\big\|_{{\mathcal{L}^{q_{0}}}(\Omega)}$ $\leq\frac{1}{6}\big\|{u}(\cdot,s)\big\|_{{\mathcal{L}^{\infty}}(\Omega)}+C_{12},$ (2.8)
where $\widehat{\sigma}$ is the dual exponent of $\sigma$ and ${\zeta_{2}=\frac{1}{q_{0}\widehat{\sigma}k}\in(0,1)}$, for all $s\in(0,T_{\mathrm{max}})$. Also using the Interpolation inequality to the last term in (2.6), we get
$\big\|{u}^{m}(\cdot,s)\big\|_{{\mathcal{L}^{q_{0}}}(\Omega)}=$ $\Big\|{u}(\cdot,s)\Big\|^{m}_{\mathcal{L}^{q_{0}m}(\Omega)}$
@@
where ${\zeta_{3}=\frac{1}{q_{0}m}\in(0,1)}$, for all $s\in(0,T_{\mathrm{max}})$ and $C_{13}>0$. Once again the Young’s inequality with $q_{0}>\frac{n}{2}$, gives
$\big\|{u}^{m}(\cdot,s)\big\|_{{\mathcal{L}^{q_{0}}}(\Omega)}\leq$ $C_{13}\Big\|{u}(\cdot,s)\Big\|_{{\mathcal{L}^{\infty}}(\Omega)}^{m-\frac{1}{q_{0}}}\leq\frac{1}{6}\big\|{u}(\cdot,s)\big\|_{{\mathcal{L}^{\infty}}(\Omega)}+C_{14},$ (2.9)
where $C_{14}>0$. Inserting (2.7)-(2.9) in (2.6), it follows that
$\sup_{t>0}\big\|{u}(\cdot,t)\big\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq\frac{1}{2}\sup_{t>0}\big\|{u}(\cdot,t)\big\|_{{\mathcal{L}^{\infty}}(\Omega)}+C_{15},$
for all $t\in(0,T_{\mathrm{max}})$, where $C_{15}>0$. Finally, we conclude that
$\big\|{u}(\cdot,t)\big\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C_{16},\qquad\qquad\qquad\forall\,t\in(0,T_{\mathrm{max}}).$ (2.10)
where the constant $C_{16}$ is positive. Initial data regularity $v_{0},w_{0}\in{\mathcal{W}^{1,q}}(\Omega)$ and the smoothing property of the heat semigroup ensure $v,w\in\mathcal{L}^{\infty}((0,T_{\mathrm{max}});{\mathcal{W}^{1,q}}(\Omega))$. This will imply that actually
$\sup_{t>0}\Big(\big\|{u}(\cdot,t)\big\|_{{\mathcal{L}^{\infty}}(\Omega)}+\big\|{v}(\cdot,t)\big\|_{{\mathcal{W}^{1,q}}(\Omega)}+\big\|{w}(\cdot,t)\big\|_{{\mathcal{W}^{1,q}}(\Omega)}\Big)<\infty.$
This completes the proof. ∎
@@ 3 Global existence of classical solutions
In this section, we prove the global existence and boundedness of the solution to (1.6). First we derive ${\mathcal{L}^{p}}(\Omega)$ norm for u{u}. Given any $s_{0}\in(0,T_{\mathrm{max}})$ such that $s_{0}<1$, Lemma 2.1 gives ${u}(\cdot,s_{0}),{v}(\cdot,s_{0}),{w}(\cdot,s_{0})\in{\mathcal{C}^{2}(\overline{\Omega})}$ with $\frac{\partial{v}(\cdot,s_{0})}{\partial\nu}=0$ and $\frac{\partial{w}(\cdot,s_{0})}{\partial\nu}=0$, we pick $C>0$ such that
$\left\{\begin{array}[]{rrll}\hskip-14.22636pt\hfil&&\sup\limits_{0\leq s\leq s_{0}}\|{u}(\cdot,s)\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C,\qquad\sup\limits_{0\leq s\leq s_{0}}\|{v}(\cdot,s)\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C,\qquad\sup\limits_{0\leq s\leq s_{0}}\|{w}(\cdot,s)\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C,\\ \\ \hskip-14.22636pt\hfil&&\|\Delta{v}(\cdot,s_{0})\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C\qquad\mathrm{and}\qquad\|\Delta{w}(\cdot,s_{0})\|_{{\mathcal{L}^{\infty}}(\Omega)}\leq C.\end{array}\right$. (3.4)
@@
Suppose that $\Omega\subset\mathbb{R}^{n}\>(n\geq 2)$, is a bounded domain with smooth boundary. Assume that $d_{1}>0,\mu>0$ and $M_{0}<1$. If $m$ satisfies
$1<m<1+\frac{2}{n},\quad n\geq 3,$
$1<m<2,\quad\qquad n=2,$
then for any $1<p<\infty$, we have
$\int_{\Omega}(1-M_{0}){{u}}^{p+m-1}$ $\leq\frac{2d_{1}(p-1)}{p^{2}\mu}\int_{\Omega}\big|{\nabla{u}}^{\frac{p}{2}}\big|^{2}+C,$
@@
Choosing $1<p^{\prime}<p+m-1$ and using the Hölder’s inequality, to get
$\int_{\Omega}{{u}}^{p+m-1}\leq$ $\left(\int_{\Omega}\left({u}^{\frac{p}{2}}\right)^{\frac{2\gamma(p+m-1)}{p}\frac{rp}{2\gamma(p+m-1)}}\right)^{\frac{2\gamma(p+m-1)}{rp}}$
$\times\left(\int_{\Omega}\left({u}^{\frac{p}{2}}\right)^{\frac{2(1-\gamma)(p+m-1)}{p}\frac{p}{2(1-\gamma)(p+m-1)}\frac{2p^{\prime}}{p}}\right)^{\frac{2(1-\gamma)(p+m-1)}{p}\frac{p}{2p^{\prime}}}$
$\leq$ $\>\Big\|{u}^{\frac{p}{2}}\Big\|^{\frac{2\gamma(p+m-1)}{p}}_{{\mathcal{L}^{r}}(\Omega)}\>\>\Big\|{u}^{\frac{p}{2}}\Big\|^{\frac{2(1-\gamma)(p+m-1)}{p}}_{{\mathcal{L}^{\frac{2p^{\prime}}{p}}}(\Omega)},$ (3.5)
where
is the exponent from the Hölder’s inequality and $r$ satisfies
$r=\left\{\begin{array}[]{rrll}\hskip-142.26378pt\hfil&&\frac{2n}{n-2},&\quad\mbox{if}\quad n\geq 3,\\ \\ \hskip-142.26378pt\hfil&&\left(\frac{2(p+m-1)}{p},\infty\right),&\quad\mbox{if}\quad n=2,\\ \\ \hskip-142.26378pt\hfil&&\infty,&\quad\mbox{if}\quad n=1.\end{array}\right$.
Using the Sobolev embedding theorem, we get the estimate from (3.5) with $C_{1}>0$
@@
where $a=\frac{2(1-\gamma)(p+m-1)}{p}\frac{1}{1-\frac{\gamma(p+m-1)}{p}}$ and $C_{2}>0$. Take $p^{\prime}>\frac{n(m-1)}{2}$ with simple computation, we ensure that
@@
Now, we estimate the terms in (3.9) as follows